The Metallica Heavy Metal Mining (MHMM) Corporation wants to diversify its operations. Some recent financial information for the company is shown here:

  Stock price $ 35
  Number of shares 25,000
  Total assets $ 8,800,000
  Total liabilities $ 2,400,000
  Net income $ 510,000

   

MHMM is considering an investment that has the same PE ratio as the firm. The cost of the investment is $700,000, and it will be financed with a new equity issue.

   

The ROE on the investment would have to be  percent (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) if we wanted the price after the offering to be $35 per share (assume the PE ratio remains constant), and the NPV of the investment would be $ (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations.). Accounting dilution (Click to select)doesdoes not occur in this case. Market value dilution (Click to select)doesdoes not occur in this case.
Answer #1

Answer:

Before new investment:

Current Equity = Total assets - Total liabilities = 8800000 - 2400000 = $6,400,000

Current EPS = Net Income / Number of shares = 510000 / 25000 = $20.40

Current PE ratio = Market price / EPS = 35 / 20.40 = 1.7156862745

Current book value per share = 6400000 / 25000 = $256

After Investment:

New Investment required = $700,000

It will be financed with a new equity issue.

Number of new shares to be issued = 700000 / 35 = 20,000

PE ratio remains constant and share price for new offering is $35

New EPS = Share price / PE ratio = 35 / 1.7156862745 = $20.40

Additional net income from new investment must have to be EPS times number of new shares issued

Additional net income = 20.40 * 20000 = $408,000

Hence:

New ROE = 408000 / 700000 = 58.29%

New ROE = 58.29%

NPV of the project = New market value of firm - Current market value of firm - Cost of the project

= $35 * (25000 + 20000) - $35 * 25000 - 700000

= $0

NPV of the project = $0

Book value per share after investment = (6400000 + 700000) / (25000 + 20000) = $157.78

Book value per share has diluted from $256 to $157.78.

Hence:

Accounting dilution does occur in this case.

Since NPV of the new project is zero, market value dilution does not occur in this case.